with a north wind blowing at 20 mph, a helicopter flies due north for 30 mi an then returns. if the total time was 48 min, find the helicopter's speed in still air.
Answer by stanbon(75887)

with a north wind blowing at 20 mph, a helicopter flies due north for 30 mi an then returns. if the total time was 48 min, find the helicopter's speed in still air.
---------------------------------
Distance = 30 mi ; time = 48/60 = 4/5 hr ; rate = distance/time = 30/(4/5)
= 150/4 = 37.5 mph
---
plane speed - wind speed = 37.5
plane speed - 20 = 37.5
plane speed in still air = 57.5 mph
=====================================
Cheers,
Stan H.